OpenStudy (anonymous):

What is the sum of the first five terms of the geometric sequence in which a1 = 5 and r = 1/5?

OpenStudy (anonymous):

I'm thinking 1/125... but idk for sure... :( is that the answer?? 1/125?? **no answer choices!!

OpenStudy (anonymous):

do u get this??? :)

OpenStudy (anonymous):

do you know the formula for geometric sequences?

OpenStudy (anonymous):



OpenStudy (anonymous):

and then i did this... a1 = 5 a2 = 5 x 1/5 = 1 a3 = 1 x 1/5 = 1/5 a4 = (1/5) x (1/5) = 1/25 a5 = (1/25) x (1/5) = 1/125 so thats how i got 1/125... idk if i did it right.. is that correct??

OpenStudy (anonymous):

@Brittni0605 @jim_thompson5910 :) :)

jimthompson5910 (jim_thompson5910):

you can use the formula, or add up those 5 terms you just found

OpenStudy (anonymous):

kk so did i do it right??? like the math part above?? so i add all them together right?? or is my answer just 1/125??

OpenStudy (anonymous):

agreed with @jim_thompson5910

OpenStudy (anonymous):

wait so i add theM??? a1 = 5 a2 = 5 x 1/5 = 1 a3 = 1 x 1/5 = 1/5 a4 = (1/5) x (1/5) = 1/25 a5 = (1/25) x (1/5) = 1/125 those 5 terms?? :D

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

so i need to convert to the same denominator right?

jimthompson5910 (jim_thompson5910):

or use the formula S = a[ (1-r^n)/(1-r) ] S = 5[ (1-(1/5)^5)/(1-1/5) ] S = ???

OpenStudy (anonymous):

S=6.224279835??? is that right??? oh and i need to write it in fraction form,... :/ forgot to say that!!